The radius of a circle is about 4 cm. The area of the circle is about what?
Answer by feliz1965(151)

Area = pir^2, where pi = 3.14 and r = radius.
A = pir^2
A = pi(4)^2
A = 16pi....This is the exact answer for the area of this circle.
If you want the approximate area answer, you must use 3.14 for pi.
A = 3.14 times 16
A = 50.24cm^2
Feliz